Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
casanueva, pinto, menzione, spinicelli, capocasa, chiummo, flaminio - 22:42 Thursday 30 June 2022 (56318) Print this report
New TCS working point exploration

The first part of the shift was dedicated to continuing with the morning activity of exploring the input beam alignment. The target was to improve the shape of the sidebands, that showed a vertical alignment mode. An offset was added to the BPC shift DOFs and we checked the behaviour of the dithering error signals and the input mirror centering error signals. A dedicated entry will be posted.

Then we tried to recover the lock of the CITF and progress with the lock acquisition. HOwever, we were not able to lock the CITF in a stable way. The few locks we had were achieved by misaligning the SR significantly (10urad). We were not able to realign the SR and keep a stable lock. Moreover, at some moments the SR local controls were showing a big movement of the SR, of several urads, without any loop being closed on it. The maximum sidebands we reached were 0.06 for the 12MHz and barely 1mW for B4 DC. Figure 1 shows a trend of the CITF lock attempts, with the SR local controls as well.

We contacted Claudia at the end of the shift and we leave the ITF in single bounce for the TCS activity.

Images attached to this report
Comments to this report:
ruggi - 23:57 Thursday 30 June 2022 (56319) Print this report

The bad behaviour of SR angular control can be better understood adding a plot of the longitudinal correction. Saturation of actuators and bad balance of actuators can cause angular oscillations when they occur together. This is a typical one year old situation, occuring when the lock acquisition strategy was far from being well tuned. An activity on SR actuators balancing could mitigate the problem of angular stability, but very likely will not fix the problem of lock acquisition.

Images attached to this comment
Search Help